# John Mendonça1,2, Isabel Trancoso1,2    Alon Lavie3,4

### v2

@@ 5.1 Benchmark Curation
In order to construct a meta-evaluation benchmark that maximizes the representation of dialogues of diverse quality, we employed a multi-stage, balanced sampling strategy to select 100 dialogues per target language.
@@
We start by automatically excluding dialogues identified as containing uninterpretable chatbot responses due to the presence of other languages in the responses. This is done with simple keyword matching of GPT-4.1 assessments. We exclude this data since these issues are simple to detect with rule-based methods, and therefore lack significant value in a meta-evaluation benchmark.
@@
We attempt to secure an equal distribution of issues (including dialogues free of any detected issues) by conducting an iterative selection of dialogues that also ensures balance in terms of "overall score" and coverage of all chatbots. This is achieved by (1) sampling a dialogue for each user-chatbot pair that contains no issues, and (2) sequentially selecting dialogues that contain each type of issue while minimizing "overall score" and chatbot representation imbalances. Additional details are found in Appendix E.
@@ 5.2 Human Annotations
The annotators recruited to conduct the evaluation of the human-likeness of user generated utterances (§3) were also tasked with evaluating the performance of the chatbot. Each annotator was asked to identify whether a given issue was present in the chatbot responses (binary classification), and in the end provide an overall assessment of the chatbot (1-5) throughout the dialogue.
We present Inter-Annotator Agreement (IAA) agreement scores in Table 1, calculated with using Krippendorff’s $\alpha$ [^8] We omit bootstrap intervals due to instability caused by the low number of positive examples in an overall small sample size (100), which leads to unreliable or undefined CI.. When looking at IAA values for issue detection, we observe that average agreement values vary greatly across languages and issue types, and that they are generally within the lower end of agreement reported by Finch et al. (2023a) for behaviour labels ($0.3\leq\alpha\leq 0.8$). The exception is for the "non-factual" (!Fac) class, which generally achieves very high correlations. This is due to the relatively objective nature of this issue [^9] Annotators were allowed to consult external references to check for ”factuality”.. Agreement is consistently lower for more subjective categories such as "repetition" (Rep), highlighting the challenges of reliably annotating these issues.
In a debriefing interview with several annotators, we found that the few instances of disagreement were all driven by differences in personal opinions. This finding highlights both the challenges of creating a balanced meta-evaluation benchmark using original responses from LLM-based chatbots, and the value of human annotation in capturing a diverse range of perspectives.
@@ 5.4 Meta-evaluation Results
Model Pearson Spearman
GPT-4.1 .4471 .4657
Gemini-2.0-flash .3463 .3563
Deepseek-V3 .2683 .2970
Deepseek-R1 .4077 .4471
Gemini-2.5-flash .3984 .4077
o3 .4532 .3984
Table 2: Overall Quality correlation results. All correlation results are significant with $p<0.01$. Bold denotes best model. Language-specific performance breakdowns are provided in Appendix E.
5.4.1 Overall Quality
Table 2 presents the correlation of the various LLMs as evaluators on our curated benchmark. Among the models evaluated, GPT-4.1 achieves the highest Spearman correlation with human ratings, and is just shy of matching the Pearson correlation performance of OpenAI’s o3. Reasoning models generally outperform their non-reasoning counterparts. Deepseek-V3 exhibits the lowest correlations of all models tested. It is important to note that these models fall short of their reported correlation on previous meta-evaluation benchmarks. [^10] For example, GPT-4 achieves .588 Pearson on Topical-Chat (Liu et al., 2023a) This highlights the value of  [Uncaptioned image] MEDAL for open-domain dialogue evaluation going forward.
5.4.2 Issue Detection
Following Finch et al. (2023b), we calculate Accuracy and F1-scores per language for positive and negative occurrences by averaging results across the double human annotations, whenever available. We present F1 results in Figure 4. Additional metrics, such as precision, recall, accuracy and predicted counts, are available in Appendix E.
Looking at F1- performance, all evaluated models achieve results comparable to human expert performance (F1 of over 0.8). This indicates that identifying issue-free dialogues is a relatively easy task for current LLMs. The F1+ results, however, reveal a significant gap between human and model performance across all failure categories. While human annotators retain strong F1+ scores across all categories, the evaluated models struggle to match this performance (with "uninterpretable" and "non-factual" as exceptions). For example, the weakest models (DeepSeek-V3 and Gemini-2.0) score F1- above 0.95 on "commonsense" and "empathy", while their F1+ scores on these two categories are below 0.20. In contrast, GPT-4.1 and o3 exhibit a more balanced performance, with notably higher F1+ values, while maintaining a strong F1-. In general, we find "empathy", "commonsense", and "relevance" issues to be the hardest to detect, with F1+ below 0.4 for even the top models. These likely require a more nuanced understanding of conversational dynamics – areas where current LLMs still struggle.
@@ 6 Findings and Contributions
    MEDAL offers native multilinguality. Instead of translating existing dialogues, MEDAL generates natively, resulting in more authentic multilingual dialogue datasets.
@@
    We conduct a comprehensive evaluation of LLMs as both open-domain chatbots and as judges, finding that: (1) there are significant cross-lingual performance differences between the primary open-access LLMs acting as chatbots; and (2) current LLMs still struggle with detection of nuanced dialogue issues such as "empathy" and "commonsense".
@@ 7 Limitations
We recognize several limitations in our generation step. First, LLMs exhibit uneven global cultural awareness, which can introduce biased or stereotypical notions when left to generate culturally diverse content (Li et al., 2024a) (though we argue this approach is preferable than simply translating from English dialogues). Second, to reduce model-specific biases, we employed several small-to-medium sized LLMs spanning different sizes and families. Our framework remains model-agnostic so that any LLM may be integrated. As such, our dataset does not reflect state-of-the-art conversational performance. Finally, we acknowledge a potential bias in our user utterances, since they must conform to the notion of quality set by our LLM-Judge.
@@
While LLMs appear to showcase breakthrough performance across numerous benchmarks, many of these are rapidly becoming either compromised due to data contamination (Sainz et al., 2023), or obsolete as models surpass their difficulty. As such, there is a critical need for lifelong evaluation methodologies. MEDAL offers a promising step in this direction, by enabling on-demand generation of novel, multilingual dialogues, therefore mitigating leakage and ensuring relevance as new LLMs emerge. However, MEDAL relies on automated assessments as a screening stage to reduce human annotation load. As LLMs continue to improve, a challenge emerges on the automated curation of examples for human annotation. Future work should look into more refined evaluation acquisition policies, such as active testing (Kossen et al., 2021; Li et al., 2024c).
@@ 8 Ethical Considerations
We targeted a compensation rate which reflects a fair wage according to local costs of living. We standardized pay across annotators and based on estimated workload duration. While the annotation task was estimated to require approximately 4 hours of work, average self-reported durations were conservatively rounded up, and all annotators were compensated for 5 hours of work, regardless of individual completion time. No personal identifiable information about the annotators was stored.
@@ Acknowledgments
This research was supported by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (Responsible.AI) and by national funds through Fundação para a Ciência e a Tecnologia (FCT) with references PRT/BD/152198/2021 and DOI: 10.54499/UIDB/50021/2020 and 10.54499/PRT/BD/152198/2021. We also want to give thanks to MS Azure services and Irving Kwong for their sponsorship.
@@ C.1 Seed Context
We utilize a seed context to minimize generation biases and augment the diversity of the generated dialogues. This context is composed of 4 different axes of diversity: Scene, Persona (including Gender), Affective State and Language. We then ask an LLM to generate the first utterance of the dialogue taking into account all this information. This composes a fallback strategy, where the LLM is free to chose one of our controlled axis as the topic of the conversation, focusing first on the scene and then on the persona and affective state as backup, or even a combination of several.
Scene.
We take inspiration from Kim et al. (2023), which show that Atomic10X scenes can be used as an effective source of social commonsense when scaling synthetic data generation. In detail, we take the same sentence expansions of Atomic10X triplets (CC-BY-4.0) as the scene description for our seed context.
@@
We sample the personas from the publicly available subset of PersonaHub (Ge et al., 2024). PersonaHub (CC-BY-NC-SA-4.0) is a collection of 1 billion diverse personas automatically curated from web data, which can be used to synthesize high-quality generations at scale.
@@
Using a simple rule-based analysis of the scenes and personas, we detect an uneven distribution of gender for Atomic 10X (54% male, 46% neutral for Atomic10X), whereas PersonaHub is purely gender-neutral. Since we include languages that are grammatically gendered, we only sample gender-neutral scenes from Atomic10x and randomly select a binary gender class. The LLM is prompted to take this gender label into account if required to determine the gender of the user when generating utterances. This attempts to minimize gender bias in user generations.
@@ C.2 Generation Details
Our framework combines open-access [^11] All models were obtained from huggingface.co. with closed-sourced LLMs accessible via an API. For open-access LLMs, we use vLLM for inference on a single node with 4xA100 80GB GPUs.
@@
We set a temperature of 0.1 and limit the generation to 64 tokens to force smaller feedback. Due to the creativity imposed by the generation parameters, we set the regeneration limit of the first User turn to 10 attempts, which was sufficient to have all seed contexts accepted by the LLM-judge. For subsequent turns, we limit regeneration attempts to 5, which achieves a 90% pass-rate. Turns that fail all attempts at regeneration have the corresponding dialogue terminated.
@@
Label Definition
Uninterpretable Outputs unclear or confusing text to the point that its meaning is difficult or impossible to determine.
Unsafe Promotes or suggests harmful actions, or is offensive or threatening .
Lacks Empathy Fails to express appropriate emotional understanding or sensitivity when the context calls for it.
Lacks Commonsense Shows poor reasoning or contradicts basic, widely known everyday knowledge.
Repetitive Unnecessarily repeats the same point, phrase, or idea in a way that adds no value.
Incoherent Contradicts itself or makes statements that do not logically follow from earlier parts of the dialogue.
Irrelevant Introduces off-topic or unrelated content that does not align with the flow or subject of the conversation.
Non Factual Provides information that is objectively false or contradicts publicly verifiable facts.
Overall Overall assessment of the dialogue.
Table 4: Definition of issue labels used for the evaluation study and meta-evaluation benchmark.
@@ C.3 Statistics
We compare our our collected dialogue dataset with frequently used datasets for meta-evaluation in Table 3. In general, our dialogue subsets obtain similar lexical diversity values (calculated using MTLD [^12] We calculate MTLD with python lexical_diversity using spacy lemmas and jieba tokens for Chinese.; McCarthy and Jarvis, 2010), with languages such as Chinese and English achieving the largest diversity (over 70), whereas German achieved the lowes diversity (45.0). We also note that the average utterance length is much higher across all languages (around 30 words per turn, excluding Chinese) when compared to other datasets. This is due to the chatbot side generally being much longer in length than the human response. Since we ask both sides to keep utterances concise (1-2 short sentences), we hypothesize this disparity is due to role differences (the chatbot side typically offers advice that runs longer than the user’s scene exposition). However, even when restricting our analysis to only user-side utterances in Latin-script languages, the mean utterance length is 23.4 words, which is still higher when compared to other datasets. This finding aligns with the well-known verbosity bias in LLMs, which tend to generate longer responses, likely as a by-product of preference optimization (Saito et al., 2023). [^13] Soda generates the dialogue as a whole, therefore utterance-level verbosity is reduced..
@@ Appendix D Automatic Dialogue Evaluation
D.1 Selected Labels
A detailed definition of the issue labels is presented in Table 4. These are adapted from Finch et al. (2023a) and Mendonça et al. (2024b) for dialogue level analysis, and have been shown to be a more adequate framework for analysing chatbots than Likert scales.
D.2 Annotation Details

### v3

@@ 5.1 Benchmark Curation
In order to construct a meta-evaluation benchmark that maximizes the diversity and representativeness of dialogue quality, we employed a multi-stage, balanced sampling strategy to select 100 dialogues per target language. With a total of 600 dialogues averaging 9 turns per dialogue, our benchmark’s size is larger than other influential dialogue-level benchmarks [^7] DSTC11 (Rodríguez-Cantelar et al., 2023) evaluates 277 dialogues, FED (Mehri and Eskenazi, 2020a) 120 dialogues and ABC-Eval (Finch et al., 2023b) 400 dialogues..
@@
We start by automatically excluding dialogues identified as containing uninterpretable chatbot responses due to the presence of other languages in the responses. These are detected via keyword matching of GPT-4.1 assessments. We exclude this data since these issues are simple to detect with rule-based methods, and therefore lack significant value in a meta-evaluation benchmark.
@@
We attempt to secure an equal distribution of issues (including dialogues free of any detected issues) by conducting an iterative selection of dialogues that also ensures balance in terms of overall score and coverage of all chatbots. To achieve this, we employed a two-stage procedure: (1) we first seeded the benchmark by randomly sampling one dialogue per user–chatbot pair that contained no issues and had an overall score of 5; and (2) we then iteratively added dialogues containing different issue types, each time selecting examples that helped balance the distribution of overall scores and chatbot representation. When multiple candidates satisfied these criteria, we chose one at random, and when no suitable example was available [^8] This was frequent when looking for dialogues from stronger chatbots with rarer issues. we ignored that requirement in the selection and continued the selection process.
@@ 5.2 Human Annotations
The annotators recruited to conduct the evaluation of the human-likeness of user generated utterances (Section 3) were also tasked with evaluating the performance of the chatbot. Each annotator was asked to identify whether a given issue was present in the chatbot responses (binary classification), and in the end provide an overall assessment of the chatbot (1-5) throughout the dialogue. Instructions emphasized the need to evaluate dialogues as if they were online chit-chat interactions.
Agreement values (reported using Krippendorff’s $\alpha$) are generally within the agreement reported by Finch et al. (2023a) for behaviour labels ($0.3\leq\alpha\leq 0.8$). Inter-annotator agreement was substantial for objective dimensions such as uninterpretable and non-factual, but lower for lacks empathy and repetitive which are more nuanced. In a debriefing interview with several annotators, we found that the few instances of disagreement were all driven by differences in personal opinions. This highlights the value of human annotation in capturing a diverse range of perspectives (Plank, 2022). Additional details regarding this annotation are presented in Appendix E.1.
@@ 5.3 Meta-Evaluation of LLMs as Automated Dialogue Evaluators
Metrics.
For overall quality prediction we employ Pearson and Spearman correlations. For issue detection, we calculate classification metrics considering the positive and negative occurrences as classes by averaging results across the double human annotations, following Finch et al. (2023b). Additional metrics, such as Pearson correlation, Precision, Recall, Accuracy and prediction counts, are available in Appendix E.2.
@@ 5.4 Meta-evaluation Results
Overall Quality Prediction.
Figure 4 presents the correlation of the various LLMs as evaluators on our curated benchmark across all languages. Among the models evaluated, o3 achieves the highest correlation with human ratings on the full benchmark (closely followed by GPT-4.1) whereas Deepseek-V3 exhibits the lowest.
Issue Detection.
Looking at F1- performance (Figure 5), all evaluated models achieve results comparable to human expert performance. This indicates that identifying issue-free dialogues is a relatively easy task for current LLMs. The F1+ results, however, reveal a significant gap between human and model performance across most failure categories. While human annotators retain strong F1+ scores across all categories, the evaluated models struggle to match this performance.
5.4.1 Discussions
Models struggle with nuanced dimensions.
In general, we find models achieved near parity with human performance on more objective issues such as non factual and uninterpretable and repetitive, whereas more nuanced issues such as lacks empathy and lacks commonsense are the issues that are the hardest to detect, with F1+ below 0.4 for even the top models. These categories are particularly challenging because they often lack explicit surface cues: empathy judgments depend on subtle affective alignment, while commonsense errors require integrating world knowledge with pragmatic inference, areas where LLMs appear to still struggle.
Reasoning models stronger than their non-reasoning counterparts.
When comparing the performance of the reasoning models against the non-reasoning models we note that the reasoning models showcase better performance on both tasks, thanks for the most part due to better recall. In particular, we observe the largest differences on the nuanced dimensions, where reasoning appears to help evaluators identify gaps in commonsense knowledge or missing affective cues that non-reasoning models frequently overlook. The exception to this is GPT-4.1, where it generally matches o3 performance.
Cross-lingual variations.
Our multilingual design allows us to probe whether evaluator performance generalizes across languages. While sample size and confounding factors such as annotator variation and uneven dialogue quality make language-wise results difficult to generalise, we nevertheless observe consistent trends (Figures 4 and 13). For example, GPT-4.1 and O3 achieve relatively stable performance across all six languages, while DeepSeek-V3 significantly underperforms in Spanish and French, indicating this model may suffer from training data imbalances or limited cross-lingual generalization for these languages.
@@ 6 Findings and Contributions
    We conduct a comprehensive evaluation of LLMs as multilingual chatbots finding that there are significant cross-lingual performance differences between the primary open-access LLMs.
@@
    We benchmark LLMs as evaluators showing that models struggle with nuanced conversational dimensions (e.g., empathy, commonsense), and that reasoning models consistently outperform non-reasoning counterparts, particularly on these harder cases.
@@ 7 Limitations
We recognize several limitations in our generation step. First, LLMs exhibit uneven global cultural awareness, which can introduce biased or stereotypical notions when left to generate culturally diverse content (Li et al., 2024a) (though we argue this approach is preferable than simply translating from English dialogues). Second, to reduce model-specific biases, we employed several small-to-medium sized LLMs spanning different sizes and families. Our framework remains model-agnostic so that any LLM may be integrated. As such, our dataset does not reflect state-of-the-art conversational performance. Finally, while our validation shows high human-likeness, future work should investigate whether similar LLM-judge failure modes appear in dialogues with genuine human users, who may exhibit more unpredictable and diverse behaviors.
@@
While LLMs appear to showcase breakthrough performance across numerous benchmarks, many of these are rapidly becoming either compromised due to data contamination (Sainz et al., 2023), or obsolete as models surpass their difficulty. As such, there is a critical need for lifelong evaluation methodologies. MEDAL offers a promising step in this direction, by enabling on-demand generation of novel, multilingual dialogues, therefore mitigating leakage and ensuring relevance as new LLMs emerge. However, MEDAL relies on automated assessments as a screening stage to reduce human annotation load. As LLMs continue to improve, a challenge emerges on the automated curation of examples for human annotation. Future work should look into more refined evaluation acquisition policies, such as active testing (Kossen et al., 2021; Li et al., 2024b).
@@ 8 Ethical Considerations
We targeted a compensation rate which reflects a fair wage according to local costs of living. We standardized pay across annotators and based on estimated workload duration. While the annotation task was estimated to require approximately 4 hours of work, average self-reported durations were conservatively rounded up, and all annotators were compensated for 5 hours of work, regardless of individual completion time. No personal identifiable information was stored.
@@ Acknowledgments
We would like to thank Bruno Martins and Luis Fernando D’Haro for their helpful feedback. This research was supported by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (Responsible.AI) and by national funds through Fundação para a Ciência e a Tecnologia (FCT) with references PRT/BD/152198/2021 and DOI: 10.54499/UIDB/50021/2020 and 10.54499/PRT/BD/152198/2021. We also want to give thanks to MS Azure services and Irving Kwong for their sponsorship.
@@ Appendix A Meta-evaluation Benchmarks
XDial-Eval (Zhang et al., 2023) is built on top of several existing English dialogue evaluation datasets, all of which predate the introduction of LLMs for dialogue. This curation process yields grand total of 14,930 annotated turns and 8,691 annotated dialogues. A commercial machine translation system is then used to extend these datasets to nine other languages: Chinese (ZH), Spanish (ES), German (DE), French (FR), Japanese (JA), Korean (KO), Hindi (HI), Arabic (AR), and Russian (RU).
@@ C.1 Seed Context
We utilize a seed context to minimize generation biases and augment the diversity of the generated dialogues. This context is composed of 4 different axes of diversity: Scene, Persona (including Gender), Affective State and Language. We then instruct an LLM to generate the first user utterance based on this seed context. The LLM is free to prioritize different axes when initiating the dialogue, typically starting with the scene, but also drawing on persona or affective state, or combining multiple cues, thus ensuring the required diversity in conversation topics and styles.
Scene description.
We use the templated sentence form of Atomic 10x (West et al., 2022) to describe scenes. Atomic 10x (CC-BY-4.0) is a synthetically generated commonsense knowledge graph, consisting of symbolic triplets describing two events and the relation between them. Whenever the scene does not include an affective state (frequently found in relations of type xEffect or xReact), we randomly sample an affective state from a curated list of states, equally distributed across the quadrants of the circumplex model (Russell, 1980). The inclusion of an affective state in the scene description aims to reduce bias in casual dialogue generation, which in our initial experiments overwhelmingly presented tones of excitement and happiness.
@@
Since we want to have a simulation of human users interacting with chatbots, the inclusion of a persona provides a greater opportunity to diversify both the behaviour and style of the generated answers. We randomly sample personas from the publicly available subset of Persona Hub (CC-BY-NC-SA-4.0) (Ge et al., 2024), a collection of diverse personas automatically curated from web data. Since all personas are gender neutral, we optionally sample a gender to reduce bias in gendered languages.
@@
Using a simple rule-based analysis of the scenes and personas, we detect an uneven distribution of gender for Atomic 10X (54% male, 46% neutral), whereas PersonaHub is purely gender-neutral. Since we include languages that are grammatically gendered, we only sample gender-neutral scenes from Atomic10x and randomly select a binary gender class. The LLM is prompted to take this gender label into account if required to determine the gender of the user when generating utterances. This attempts to minimize gender bias in user generations.
First Turn Generation.
@@ C.2 Generation Details
Our framework combines open-access [^9] All models were obtained from huggingface.co. with closed-sourced LLMs accessible via an API. For open-access LLMs, we use vLLM for inference on a single node with 4xA100 80GB GPUs. Dialogue length is capped at 10 turns due to cost constraints.
@@
We set a temperature of 0.1 and limit the generation to 64 tokens to force smaller feedback. Due to the creativity imposed by the generation parameters, we set the regeneration limit of the first User turn to 10 attempts, which was sufficient to have all seed contexts accepted by the LLM-judge. For subsequent turns, we limit regeneration attempts to 5, which achieves a 90% pass-rate, and keeps costs below a factor of 4101010Incorporating the feedback loop makes the process 4 times more expensive in terms of LLM queries, assuming response generation and evaluation steps have equal cost.. Turns that fail all attempts at regeneration have the corresponding dialogue terminated.
@@ C.3 Statistics
We compare our our collected dialogue dataset with frequently used datasets for meta-evaluation in Table 2. In general, our dialogue subsets obtain similar lexical diversity values (calculated using MTLD [^11] We calculate MTLD with python lexical_diversity using spacy lemmas and jieba tokens for Chinese.; McCarthy and Jarvis, 2010), with languages such as Chinese and English achieving the largest diversity (over 70), whereas German achieved the lower diversity (45.0). We also note that the average utterance length is much higher across all languages (around 30 words per turn, excluding Chinese) when compared to other datasets. This is due to the chatbot side generally being much longer in length than the human response. Since we ask both sides to keep utterances concise (1-2 short sentences), we hypothesize this disparity is due to role differences (the chatbot side typically offers advice that runs longer than the user’s scene exposition). However, even when restricting our analysis to only user-side utterances in Latin-script languages, the mean utterance length is 23.4 words, which is still higher when compared to other datasets. This finding aligns with the well-known verbosity bias in LLMs, which tend to generate longer responses, likely as a by-product of preference optimization (Saito et al., 2023) [^12] Soda generates the dialogue as a whole, therefore utterance-level verbosity is reduced..
C.4 Translation vs Native Generation
To validate our hypothesis that natively generated dialogues in a target language are more authentic than to translated ones, we conduct a head-to-head comparison. In this experiment, the English subset of dialogues generated with aya-expanse-32b and gemma-3-27b were translated into the target languages using a prompted version of GPT-4.1. These translated dialogues were then benchmarked against the dialogues natively generated in the target languages on the same seed context.
We use gemini-2.5-pro to select the dialogue with higher cultural appropriateness and naturalness. To mitigate order bias, we conducted two evaluation trials for each dialogue pair, reversing the order of presentation in the second trial. A "win" for native generation occurs when gemini-2.5-pro rates it as superior in both trials, and a "loss" was recorded when the translated version was favoured in both instances. All other outcomes were classified as a "tie".
Figure 7: Automated evaluation results from a head-to-head comparison between natively generated and translated dialogues. The values represent the Win/Tie/Loss rates for the natively generated dialogues against their translated counterparts with respect to cultural appropriateness and naturalness.
The results presented in Figure 7 indicate that for most of the languages tested, dialogues generated natively are better in cultural appropriateness and naturalness compared to those translated from English. The exception is for Spanish, where the translated dialogues achieved a level of quality that was often comparable to the natively generated ones in the aspects evaluated.
@@ Appendix D Automatic Dialogue Evaluation
D.1 Annotation Details